OpenStudy (anonymous):

Simplify 6 4^√4-7 3^√4+2 3^√4-3 4^√4

OpenStudy (anonymous):

\[6\sqrt[4]{4-7}\sqrt[3]{4+2}\sqrt[3]{4-3}\sqrt[4]{4}\]

OpenStudy (anonymous):

multiply \[x^{1/3} \times x^{3/7}\]

OpenStudy (anonymous):

I need help with these two @AriPotta

OpenStudy (anonymous):

1.1760

OpenStudy (anonymous):

thats not even close,...

OpenStudy (aripotta):

i can help with that second one, but idk about that first one when you multiply exponents, you add them 1/3 + 3/7 = 7/21 + 9/21 = 16/21 so x^(16/21)

OpenStudy (anonymous):

alright i think i got the first one now

OpenStudy (anonymous):

thank (: and so whatd you get

OpenStudy (anonymous):

-713

OpenStudy (anonymous):

\[64^{2}-73^{2}+23^{2}-3^{2}\]

OpenStudy (aripotta):

oh ok! you should have posted that to begin with lol

OpenStudy (aripotta):

combine your like terms

OpenStudy (anonymous):

i dont have my calculator that was the final step do the exponents then subtract and add

OpenStudy (anonymous):

how ?

OpenStudy (aripotta):